OpenStudy (anonymous):

A manufacturer wants to design an open box (no top) having a square base and a surface area of 300 square inches. What dimensions will produce a box with maximum volume?

OpenStudy (anonymous):

I don't even know where to start on this. :(

OpenStudy (anonymous):

thats not right. I know i have to maximize it but just don't know how.

OpenStudy (anonymous):

it is using derivative

OpenStudy (anonymous):

normally, if the question wasn't so abstract I'd get it. I would probably find Surface area, then take derivative and plug in one of my variable to eliminate then do a test to see if max. I just feel like its missing some numbers, but I know its not because I know there is an answer to it...kinda remember seeing something like this in class, but didn't get it.

OpenStudy (anonymous):

I'm use to already being given the equation lol. but this one doesn't give it to me.

OpenStudy (anonymous):

the equation is: surface area = 300 = b^2+4bh

OpenStudy (kropot72):

The answer for the box dimension to give maximum volume and a surface area of 300 sq. ins. is : Height = 5 inches Width = 10 inches Depth = 10 inches Posting of solution method to follow.

OpenStudy (anonymous):

patiently waiting.

OpenStudy (anonymous):

surface area of the box is= b^2+4bh=300. volume of box is V=b^2h. h=V/b^2 => S=b^2+ 4V/b=300. => V=(300b-b^3)/4. => dV=(300-3b^2)/4=0 => b=10. substitute this in surface area equation to get height to be 5 inches.

OpenStudy (kropot72):

Let base side = b inches Let height = h Area of sides = 300 - b^2 300 - b^2 = 4 * b * h \[h=\frac{300-b ^{2}}{4b}\] \[Volume=b ^{2}(\frac{300-b ^{2}}{4b})\] \[V=75b-\frac{b ^{3}}{4}\] \[\frac{dV}{db}=75-\frac{3b ^{2}}{4}\] Putting f'(b) = 0 and solving gives b = 10 Substitution gives h = 5

OpenStudy (anonymous):

Thank you everyone for all your help. I understand now. I wish I could give out more than 1 medal because you all contributed to helping me find the answer.